Ouiz Note: It is recommended that you save your response as you complete each question. Question 1 (6 points) Huitron inc. expe its sales in September to be s 143,000. The companys contribution margin ratio is 65% and its fixed monthly expenses are $62,000. Required: at Goyer
Answer #1

Net contribution = sales * contribution ratio

= 143000*65% = 92950-62000 = $30,950.

Looking forward to your upvote.

Net operating income = 92950-